OpenStudy (anonymous):

Select all of the following that are ordered pairs of the given function. f(x) = 3 - 2x (-2, -1) (-1, 5) (0, 3) (1, 0) (2, -1)

OpenStudy (anonymous):

f(x) is the y point y=3-2x so put x=0, you will get y=3 so the pair is (0,3) so try the rest of them take the value of x then put it into the equation and see if it gives the the other number of the pair.

OpenStudy (anonymous):

do u understand? @haahahahahahahaaaa

OpenStudy (anonymous):

f(x) is the y point y=3-2x so put x=-2, you will get y=-1 so the pair is (-2,-1) so try the rest of them take the value of x then put it into the equation and see if it gives the the other number of the pair.